OpenStudy (anonymous):

Completely factor the following expression: 64x² – 176xy + 121y²

OpenStudy (anonymous):

64x^2 - 176xy + 121y^2 = 64x^2 - 88xy - 88xy +121y^2 = 8x(8x - 11y) - 11y(8x - 11y) = (8x - 11y)(8x - 11y) = (8x - 11y)^2